# Adaptive Batch Size for Privately Finding Second-Order Stationary Points

### v1

@@ 3 SOSP
Under Assumption 3.1 and with gradient oracles such that ~$\|\widetilde{\nabla}_{t}-\nabla F(x_{t})\|\leq\gamma$ for any $t\in[T]$, setting $T=\tilde{O}(B/\eta\gamma^{2})$ and $\zeta=\gamma/\sqrt{\log(d/\iota)}$ and supposing it does not halt before completing all $T$ iterations, with probability at least $1-T\iota$, at least one point in the output set $\{x_{1},\cdots,x_{T}\}$ of Algorithm 2 is $\tilde{O}(\gamma)$-SOSP.
@@
The proof intuition of Proposition 3.4 is, if we do not find an $O(\gamma)$-SOSP, then on average, the function value will at least decrease by $\Omega(\eta/\gamma^{2})$. As we know $F_{\mathcal{P}}(x_{0})\leq F_{\mathcal{P}}^{*}+B$, hence $O(B/\eta\gamma^{2})$ steps can ensure we find an $O(\gamma)$-SOSP. See more discussion on Proposition 3.4 in the Appendix.
Algorithm 2 Stochastic Spider
1:Input: Dataset $\mathcal{D}$, privacy parameters $\varepsilon,\delta$, parameters of objective function $B,M,G,\rho$, parameter $\kappa$, failure probability $\omega$, batch size parameter $b$, noise parameter $\zeta$
2:set $\mathrm{drift}_{0}=\kappa,\mathrm{frozen}_{-1}=1,\Delta_{-1}=0,\mathcal{D}_{r}\leftarrow\mathcal{D},t=0$
3:while $t<T$ and the number of unused functions is larger than $b$ do
4:     if $\mathrm{drift}_{t}\geq\kappa$ then
5:         $\nabla_{t}=\mathcal{O}_{1}(x_{t})$, $\mathrm{drift}_{t}=0$, $\mathrm{frozen}_{t}=\mathrm{frozen}_{t-1}-1$
6:     else
7:         $\Delta_{t}=\mathcal{O}_{2}(x_{t},x_{t-1})$, $\nabla_{t}=\nabla_{t-1}+\Delta_{t}$, $\mathrm{frozen}_{t}=\mathrm{frozen}_{t-1}-1$
8:     end if
9:     if ~$\|\widetilde{\nabla}_{t-1}\|\leq\gamma\log^{3}(BMd/\rho\omega)\bigwedge\mathrm{frozen}_{t-1}\leq 0$ then
10:         Set $\mathrm{frozen}_{t}=\Gamma,\mathrm{drift}_{t}=0$
11:         Set ~$\widetilde{\nabla}_{t}=\nabla_{t}+\mathrm{TREE}(t)+g_{t}$, where $g_{t}\sim\mathcal{N}(0,\frac{\zeta^{2}}{d}I_{d})$
12:     else
13:         Set ~$\widetilde{\nabla}_{t}=\nabla_{t}+\mathrm{TREE}(t)$
14:     end if
15:     $x_{t+1}=x_{t}-\eta\widetilde{\nabla}_{t},\mathrm{drift}_{t+1}=\mathrm{drift}_{t}+\|\widetilde{\nabla}_{t}\|_{2}$, $t=t+1$
16:end while
17:Return: $\{x_{1},\cdots,x_{t}\}$
Algorithm 2 follows the SpiderBoost framework. We either query $\mathcal{O}_{1}$ to estimate the gradient itself or query $\mathcal{O}_{2}$ to estimate the gradient difference over the last consecutive points. The term $\mathrm{drift}$ controls the estimated error. When $\mathrm{drift}_{t}$ is small, we know $\Delta_{t}$ is still a good estimator, and when $\mathrm{drift}_{t}$ is large, we draw a fresh estimator from $\mathcal{O}_{1}$. The term $\mathrm{frozen}$ is used for the technical purpose of applying Lemma 3.3. When we meet a potential saddle point, we add Gaussian noise $g_{t}$ to escape from the saddle point and set $\mathrm{frozen}$ to be $\Gamma$; this ensures that we won’t add the Gaussian again in the following $\Gamma$ steps.
3.1 Constructing Private Gradient Oracles
@@
We construct the gradient oracles below in Algorithm 3.
Algorithm 3 gradient oracles
1:gradient $\mathcal{O}_{1}$
@@
1:gradient $\mathcal{O}_{2}$
@@
Lemma 3.5 (Gradient oracles with bounded error).
$\Pr[\|\mathcal{O}_{1}(x_{t})-\nabla F(x_{t})\|\geq\zeta_{1}]\leq\iota,$
where $\zeta_{1}=O((G\sqrt{\log(d/\iota)/b})$.
where $\zeta_{2}=O(M\|x_{t}-x_{t-1}\|\sqrt{\log(d/\iota)/b_{t}})$.
@@
For each data $z\sim\mathcal{P}$, we know
@@
Then the conclusion follows from Lemma 2.4.
@@
From now on, we adopt Algorithm 3 as the gradient oracles for Line 5 and Line 7 respectively in Algorithm 2, and we set $\eta=1/M$. We then bound the error between gradient estimator $\nabla_{t}$ and the true gradient $\nabla F(x_{t})$ for Algorithm 2.
Lemma 3.6.
$\|\nabla_{t}-\nabla F(x_{t})\|^{2}\leq C(\frac{G^{2}}{b}+\sum_{i=\tau_{t}+1}^{t}M^{2}\|x_{i}-x_{i-1}\|^{2}/b_{i})\log(Td/\iota).$ (1)
@@
That is $\Delta_{i}-(\nabla F(x_{i})-\nabla F(x_{i-1}))$ is zero-mean and $\mathrm{nSG}(M\|x_{i}-x_{i-1}\|\sqrt{\log(dT/\iota)}/\sqrt{b_{i}})$ by applying Lemma 3.5. Then Equation (1) follows from applying Lemma 2.4. ∎
Now, we consider the noise added from the tree mechanism to make the algorithm private.
Lemma 3.7.
If we set $\sigma=(\frac{G}{b}+\max_{t}\frac{\|\widetilde{\nabla}_{t}\|}{b_{t}})\log(1/\delta)/\varepsilon$, in the tree mechanism (Algorithm 1) and use Algorithm 3 as gradient oracles, then Algorithm 2 is $(\varepsilon,\delta)$-DP.
@@
It suffices to consider the sensitivity of the gradient oracles.
@@
The privacy guarantee follows from the tree mechanism (Theorem 2.2). ∎
With the noise added from the tree mechanism in mind, now we get the high-probability error bound of our gradient estimators ~$\widetilde{\nabla}_{t}$.
Lemma 3.8.
In Algorithm 2, setting $b=G\sqrt{d}/\varepsilon\alpha+G^{2}/\alpha^{2},b_{t}=\max\{\frac{\|\widetilde{\nabla}_{t}\|\cdot\sqrt{d}}{\alpha\varepsilon},\frac{\kappa\cdot\|\widetilde{\nabla}_{t}\|}{\alpha^{2}},1\}$, and $\sigma=2\log(1/\delta)\alpha/\sqrt{d}$ correspondingly according to Lemma 3.7, for each $t\in[T]$, we know Algorithm 2 is $(\varepsilon,\delta)$-DP, and with probability at least $1-\iota$, ~$\|\widetilde{\nabla}_{t}-\nabla F(x_{t})\|\leq\gamma$, where $\gamma=\tilde{O}(\alpha).$
@@
By our setting of parameters, we know
$(\frac{G}{b}+\max_{t}\frac{\|\widetilde{\nabla}_{t}\|}{b_{t}})\leq 2\varepsilon\alpha/\sqrt{d}.$
Then our choice of $\sigma$ ensures the privacy guarantee by Lemma 3.7.
For any $t\in[T]$, we have
@@
$\|\nabla_{t}-\nabla F(x_{t})\|^{2}\lesssim$ $(\alpha^{2}+\sum_{i=\tau_{t}+1}^{t}\|\widetilde{\nabla}_{t}\|^{2}/b_{t})\log(nd/\iota)$
$\lesssim$ $(\alpha^{2}+\sum_{i=\tau_{t}+1}^{t}\|\widetilde{\nabla}_{t}\|\cdot\min\{\frac{\alpha\varepsilon}{\sqrt{d}},\alpha^{2}/\kappa\})\log(nd/\iota)$
@@
Hence we conclude that $(2)\lesssim\alpha\sqrt{\log(nd/\iota)}$, which completes the proof. ∎
Lemma 3.9.
$\tilde{O}\Big{(}\frac{bBM}{\kappa\gamma}+BM(\frac{\sqrt{d}}{\gamma^{2}\varepsilon}+\frac{\kappa}{\gamma^{3}}+\frac{1}{\gamma^{2}})\Big{)}.$

### v2

@@ 3.2 Main Algorithm
Under Assumption 3.1 and with oracles such that ~$\|\widetilde{\nabla}_{t}-\nabla F(x_{t})\|\leq\gamma$ and $\|H_{t}-\nabla^{2}F(x_{t})\|_{2}\leq\sqrt{\rho\gamma}/2$ for any $t\in[T]$. When $\sigma_{\mathrm{tree}}=\sigma_{t}=\frac{\gamma}{\sqrt{d\log(T/\iota)}}$, $\Gamma=\tilde{O}(M/\sqrt{\rho\gamma})$, $\varkappa\Xi\leq\gamma,\eta=1/M$ and $\Xi=\sqrt{\gamma/\rho}$, setting $T=\tilde{O}(B/\eta\gamma^{2})$, and supposing it does not halt before completing all $T$ iterations, with probability at least $1-T\iota$, at least one point in the output set $\{x_{1},\cdots,x_{T}\}$ of Algorithm 3 is $\tilde{O}(\gamma)$-SOSP.
@@
The proof intuition of Proposition 3.10 is, if we do not find an $O(\gamma)$-SOSP, then on average, the function value will at least decrease by $\Omega(\eta/\gamma^{2})$. As we know $F_{\mathcal{P}}(x_{0})\leq F_{\mathcal{P}}^{*}+B$, hence $O(B/\eta\gamma^{2})$ steps can ensure we find an $O(\gamma)$-SOSP. See the proof of Proposition 3.10 in the Appendix.
@@
It suffices to build the private oracles with provable utility guarantees. We construct the gradient oracles below in Algorithm 4.
Lemma 3.11 (Oracles with bounded error).
$\Pr[\|\mathcal{O}_{1}(x_{t})-\nabla F(x_{t})\|\geq\zeta_{1}]\leq\iota,$
where $\zeta_{1}=O(G\sqrt{\log(d/\iota)/b})$.
where $\zeta_{2}=O(M\|x_{t}-x_{t-1}\|\sqrt{\log(d/\iota)/b_{t}})$.
(3) for any $x_{t}$. we have $\operatorname*{\mathbb{E}}[\mathcal{O}_{3}(x_{t})]=\nabla^{2}F(x_{t})$ and
$\Pr[\|\mathcal{O}_{3}(x_{t})-\nabla^{2}F(x_{t})\|\geq\zeta_{3}]\leq\iota,$
where $\zeta_{3}=O(M\sqrt{\log(d/\iota)/b})$.
(4)for any $x_{t},x_{t-1}$, we have $\operatorname*{\mathbb{E}}[\mathcal{O}_{4}(x_{t},x_{t-1})]=\nabla^{2}F(x_{t})-\nabla^{2}F(x_{t-1})$ and
$\Pr[\|\mathcal{O}_{4}(x_{t},x_{t-1})-(\nabla F(x_{t})-\nabla F(x_{t-1}))\|\geq\zeta_{4}]\leq\iota,$
where $\zeta_{4}=O(\rho\|x_{t}-x_{t-1}\|\sqrt{\log(d/\iota)/b_{t}})$.
From now on, we adopt Algorithm 4 as the gradient oracles for Line 6 and Line 10 respectively in Algorithm 3, and we set $\eta=1/M$. We then bound the error between gradient estimator $\nabla_{t}$ and the true gradient $\nabla F(x_{t})$ for Algorithm 3.
Lemma 3.12.
$\|\nabla_{t}-\nabla F(x_{t})\|^{2}\leq C(\frac{G^{2}}{b}+\sum_{i=\tau_{t}+1}^{t}M^{2}\|x_{i}-x_{i-1}\|^{2}/b_{i})\log(Td/\iota),$ (2)
$\|H_{t}-\nabla^{2}F(x_{t})\|^{2}\leq C(\frac{M^{2}}{b}+\sum_{i=\tau_{t}+1}^{t}\rho^{2}\|x_{i}-x_{i-1}\|^{2}/b_{i})\log^{2}(Td/\iota).$ (3)
Now, we consider the noise added from the tree mechanism and the noise added in the stochastic power method to make the algorithm private.
Lemma 3.13.
If we set $\sigma_{\mathrm{tree}}=(\frac{G}{b}+\max_{t}\frac{\|\widetilde{\nabla}_{t}\|}{b_{t}})\log(1/\delta)/\varepsilon$ in the tree mechanism (Algorithm 1), $\sigma_{t}=(\frac{M}{b}+\max_{t}\frac{\rho\|x_{t}-x_{t-1}\|}{b_{t}})\cdot\frac{2\Xi\sqrt{\Gamma\tau\log(1/\delta)}}{\varepsilon}$ and use Algorithm 4 as oracles, then Algorithm 3 is $(\varepsilon,\delta)$-DP.
Algorithm 4 oracles
1:gradient oracle $\mathcal{O}_{1}$
@@
1:gradient difference oracle $\mathcal{O}_{2}$
@@
1:Hessian oracle $\mathcal{O}_{3}$
2:inputs: $x_{t}$
3:draw a batch size of $b$ among unused functions
4:return: $\frac{1}{b}\sum_{z}\nabla^{2}f(x_{t};z)$
1:Hessian difference oracle $\mathcal{O}_{4}$
2:inputs: $x_{t},x_{t-1}$
3:draw a batch size of $b_{t}$ among unused functions
4:return: $\frac{1}{b_{t}}\sum_{z}(\nabla^{2}f(x_{t};z)-\nabla^{2}f(x_{t-1};z))$
With the noise added in mind, we get the high-probability error bound of gradient estimators ~$\widetilde{\nabla}_{t}$ and Hessian estimators $H_{t}$.
Lemma 3.14.
In Algorithm 3, setting fixed batch size $b=G\sqrt{d}/\varepsilon\alpha+G^{2}/\alpha^{2}+\frac{M\sqrt{d}}{\rho\varepsilon}+\frac{M^{2}}{\rho\alpha}$, adaptive batch size $b_{t}=\max\{\frac{\|g_{t}\|\cdot\sqrt{d}}{\alpha\varepsilon},\frac{\kappa\cdot\|g_{t}\|}{\alpha^{2}},\frac{\rho\kappa\cdot\|g_{t}\|}{M^{2}\alpha},1\}$, $\Xi=\sqrt{\gamma/\rho},\tau=\frac{\alpha^{3/2}}{M\sqrt{\rho}},\Gamma=\tilde{O}(M/\sqrt{\rho\gamma})$, $\sigma_{t}=2\log(1/\delta)\alpha/\sqrt{d},\forall t$ and $\sigma_{\mathrm{tree}}=2\log(1/\delta)\alpha/\sqrt{d}$ correspondingly according to Lemma 3.13, for each $t\in[T]$, we know Algorithm 3 is $(\varepsilon,\delta)$-DP, and with probability at least $1-\iota$, $\|g_{t}-\nabla F(x_{t})\|\leq\gamma$, $\|H_{t}-\nabla^{2}F(x_{t})\|_{2}\leq\sqrt{\rho\gamma}/2$, where $\gamma=\tilde{O}(\alpha).$
Lemma 3.15.
$\tilde{O}\Big{(}\frac{bBM}{\kappa\gamma}+BM(\frac{\sqrt{d}}{\gamma^{2}\varepsilon}+\frac{\kappa}{\gamma^{3}}+\frac{\rho\kappa}{M^{2}\gamma^{2}}+\frac{1}{\gamma^{2}})\Big{)}.$
Lemma 3.16.
$\alpha=$ $O\Big{(}((BGM)^{1/3}+\sqrt{BM})(\frac{\sqrt{d}}{n\varepsilon})^{1/2}+\frac{B^{\frac{2}{9}}M^{\frac{2}{9}}G^{\frac{5}{9}}+B^{\frac{4}{9}}M^{\frac{4}{9}}G^{\frac{1}{9}}}{n^{1/3}}\Big{)}$
$+O\Big{(}\frac{(MB)^{1/2}}{\sqrt{n}}+\frac{\sqrt{BM^{2}}}{\rho\sqrt{n}}+\frac{B^{1/3}M^{4/3}}{G^{1/6}\sqrt{\rho n}}+\frac{B^{2/3}G^{1/6}\sqrt{\rho}}{M^{1/3}\sqrt{n}}+\frac{B\rho\sqrt{d}}{Mn\varepsilon}\Big{)},$
Theorem 3.17.
Given $\delta>0,\varepsilon\leq O(1)$, with gradient oracles in Algorithm 4, setting $b=G\sqrt{d}/\varepsilon\alpha+G^{2}/\alpha^{2}+\frac{M\sqrt{d}}{\rho\varepsilon}+\frac{M^{2}}{\rho\alpha},b_{t}=\max\{\frac{\|g_{t}\|\cdot\sqrt{d}}{\alpha\varepsilon},\frac{\kappa\cdot\|g_{t}\|}{\alpha^{2}},\frac{\rho\kappa\cdot\|g_{t}\|}{M^{2}\alpha},1\}$, $\kappa=\max\{\frac{\alpha\sqrt{d}}{\varepsilon},(BGM)^{1/3}\}$, $\Xi=\sqrt{\gamma/\rho},\tau=\frac{\alpha^{3/2}}{M\sqrt{\rho}},\Gamma=\tilde{O}(M/\sqrt{\rho\gamma})$, and $\sigma_{\mathrm{tree}}=\sigma_{t}=\alpha/\sqrt{d},\forall t$, Algorithm 3 is $(\varepsilon,\delta)$-DP, and if the dataset is i.i.d. drawn from the underlying distribution $\mathcal{P}$, at least one of its outputed points is $\tilde{O}(\alpha)$-SOSP, where
$\alpha=$ $O\Big{(}((BGM)^{1/3}+\sqrt{BM})(\frac{\sqrt{d}}{n\varepsilon})^{1/2}+\frac{B^{\frac{2}{9}}M^{\frac{2}{9}}G^{\frac{5}{9}}+B^{\frac{4}{9}}M^{\frac{4}{9}}G^{\frac{1}{9}}}{n^{1/3}}\Big{)}$
$+O\Big{(}\frac{(MB)^{1/2}}{\sqrt{n}}+\frac{\sqrt{BM^{2}}}{\rho\sqrt{n}}+\frac{B^{1/3}M^{4/3}}{G^{1/6}\sqrt{\rho n}}+\frac{B^{2/3}G^{1/6}\sqrt{\rho}}{M^{1/3}\sqrt{n}}+\frac{B\rho\sqrt{d}}{Mn\varepsilon}\Big{)},$
Remark 3.18.
If we treat the parameters $B,G,M,\rho$ as constants $O(1)$, then we get $\alpha=\tilde{O}((\frac{\sqrt{d}}{n\varepsilon})^{1/2}+\frac{1}{n^{1/3}})$ as claimed before in the abstract and introduction.
If we make further assumptions, like assuming the functions are defined over a constraint domain $\mathcal{X}\subset\mathbb{R}^{d}$ of diameter $R$ and we allow exponential running time, we can get some other standard bounds that can be better than Theorem 3.17 in some regimes. See Appendix C for more discussions.
4 Discussion
Appendix A Omitted Proof of Subsection 3.1
A.1 Proof of Lemma 3.5
@@ 4 Discussion
By the concentration of the Gaussian, we know with probability at least $21-\iota/$ 2, $\|\zeta_{t}\|_{2}\leq\gamma$ for all $t\in[\Gamma]$. The following proof is conditional on the event that $\|\zeta_{t}\|_{2}\leq\gamma$ for all $t$.
By the Assumption 3.1, for each $t\in[\Gamma]$, we know
$\|g_{t-1}-\nabla F(x_{t-1})\|_{2}=$ $\|g+H(x_{t-1}-x_{0})+\zeta_{t-1}-\nabla F(x_{t-1})\|_{2}$
$\leq$ $\|g+H(x_{t-1}-x_{0})-\nabla F(x_{t-1})\|_{2}+\|\zeta_{t-1}\|_{2}$
$\leq$ $\|g+H(x_{t-1}-x_{0})-\nabla F(x_{t-1})\|_{2}+\gamma$
$=$ $\|g-\nabla F(x_{0})+H(x_{t-1}-x_{0})-(\nabla F(x_{t-1})-\nabla F(x_{0})\|_{2}+\gamma$
$\leq$ $2\gamma+\|(H-\nabla^{2}F(z))(x_{t-1}-x_{0})\|_{2},$
where $z$ is a point in the section between $x_{t-1}$ and $x_{0}$. Note that $\|x_{t-1}-x_{0}\|_{2}<\Xi$ by the algorithm design, hence we know $\|H-\nabla^{2}F(z)\|_{2}\leq\varkappa+\rho\Xi$. Hence, we have
$\|g_{t-1}-\nabla F(x_{t-1})\|_{2}\leq 2\gamma+(\varkappa+\rho\Xi)\Xi.$
This completes the proof.
A.2 Proof of Lemma 3.8
Proof.
By Claim 3.6, we have
$F(x_{t})-F(x_{0})\leq-\frac{\eta}{4}\sum_{i=1}^{t}\|\nabla F(x_{i})\|_{2}^{2}+5\eta\sum_{i}^{t}\|\varrho_{i}\|_{2}^{2}.$ (4)
Note that $x_{t}-x_{0}=\eta(\sum_{i=1}^{t}\nabla F(x_{i})-\varrho_{i})$, which means $\eta\|(\sum_{i=1}^{t}\nabla F(x_{i})-\varrho_{i})\|_{2}\geq\Xi$. By Lemma 3.5 and the preconditions, we know
$\|\varrho_{i}\|\leq 3\gamma.$
Hence we know
$\eta\|\sum_{i=1}^{t}\nabla F(x_{i})\|\geq\Xi-3\gamma\Gamma\geq\sqrt{\gamma/2\rho}.$
Then by Equation 4, we know
$F(x_{t})-F(x_{0})\leq$ $-\frac{\eta}{4\Gamma}(\sum_{i=1}^{t}\|\nabla F(x_{i})\|)^{2}+45\eta\Gamma\gamma^{2}$
$\leq$ $-\frac{\eta\gamma}{8\Gamma\rho}+45\eta\Gamma\gamma^{2}$
$\leq$ $-\Phi.$
A.3 Proof of Lemma 3.9
Proof.
For proof purposes, let $\{x_{t}\}_{t\in[\Gamma]}$ and $\{x_{t}^{\prime}\}_{t\in[\Gamma]}$ be the two trajectories with $x_{1}=w$ and $x_{1}^{\prime}=u$.
It suffices to show that
$\max\{\|x_{\Gamma}-x_{0}\|_{2},\|x_{\Gamma}^{\prime}-x_{0}\|_{2}\}\geq\Xi.$ (5)
Let $z_{t}=x_{t}-x_{t}^{\prime}$ be the difference. Hence we have
$z_{t+1}=$ $z_{t}-\eta[g_{t-1}-g_{t-1}^{\prime}]$
$=$ $z_{t}-\eta H(x_{t}-x_{t}^{\prime})$
$=$ $(I-\eta H)z_{t}.$
Recall that $z_{1}=\mu r\mathbf{e}_{1}$, $\lambda_{\min}(\nabla^{2}F(x_{0}))\leq-\sqrt{\rho\alpha}$ and $\lambda_{\min}(H)\leq\sqrt{\rho\alpha}/2$. Then we know that
$\|z_{t}\|\geq(1+\eta\sqrt{\rho\gamma}/2)^{t}\mu r,$
which means
$\|z_{\Gamma}\|_{2}\geq 2\Xi$
by our choices of parameters. This establishes Equation (5) and completes the proof.
A.4 Proof of Proposition 3.10
Proof of Proposition 3.10.
By Lemma 3.2 and the precondition that ~$\|\widetilde{\nabla}_{t}-\nabla F(x_{t})\|\leq\gamma$, we know that, if $\|\nabla F(x_{t})\|\geq 4\gamma$, then $F(x_{t+1})-F(x_{t})\leq-\eta\|\widetilde{\nabla}\|^{2}/16$. Otherwise, $\|\nabla F(x_{t})\|<4\gamma$. If $\|\nabla F(x_{t})\|<4\gamma$ but $x_{t}$ is a saddle point, then by Theorem 3.4, we know with probability at least $1-\iota$,
where $\Gamma=\tilde{O}(M/\sqrt{\rho\gamma})$. Then if none of the points in $\{x_{i}\}_{i\in[T]}$ is an $\tilde{O}(\gamma)$-SOSP, then we know $F(x_{T})-F(x_{0})<-B$, which is contradictory to Assumption 3.1. Hence at least one point in $\{x_{i}\}_{i\in[T]}$ should be an $\tilde{O}(\gamma)$-SOSP, and hence complete the proof.
A.5 Proof of Theorem 3.4
Now we complete the proof of Theorem 3.4.
Proof of Theorem 3.4.
By Lemma 3.8 and the definition of $\mathcal{X}(x_{0})$ (see Equation (1)), we have
$\Pr[F(y)-F(x_{0})\leq-\Phi]\geq\Pr\Big{[}x_{1}\notin\mathcal{X}(x_{0})\mid\|\zeta_{t}\|_{2}\leq\gamma,\forall t\Big{]}+\iota/2.$
It suffices to upper bound $\Pr\Big{[}x_{1}\in\mathcal{X}(x_{0})\mid\|\zeta_{t}\|_{2}\leq\gamma,\forall t\Big{]}$. Let $\mu_{0}=\iota^{2}/4\sqrt{d}$. Recall the Gaussian we add is $\sigma_{0}=\frac{\gamma}{\sqrt{d\log(T/\iota)}}$. Suppose we add two independent Gaussians $\zeta_{0},\zeta_{0}^{\prime}$ and get two independent first iterates $x_{1}$ and $x_{1}^{\prime}$. By the property of Gaussians, we have
$\Pr[|\mathbf{e}_{1},\zeta_{0}-\zeta_{0}^{\prime}|\leq\mu\gamma]=2\mathrm{erf}(\mu\gamma/\sigma_{0})\leq\iota^{2}/16.$
Then
$\Pr\big{[}|\mathbf{e}_{1},\zeta_{0}-\zeta_{0}^{\prime}|\leq\mu\gamma\mid\|\zeta_{0}\|\leq\gamma,\|\zeta_{0}^{\prime}\|\leq\gamma\big{]}\leq\frac{\iota^{2}/16}{1-\iota}\leq\iota^{2}/4.$
By Lemma 3.9 and independence between $x_{1},x_{1}^{\prime}$, we have
$\Pr\Big{[}x_{1}\in\mathcal{X}(x_{0})\mid\|\zeta_{t}\|_{2}\leq\gamma,\forall t\Big{]}\leq$ $\sqrt{\Pr[x_{1},x_{1}^{\prime}\in\mathcal{X}(x_{0})\mid\|\zeta_{0}\|\leq\gamma,\|\zeta_{0}^{\prime}\|\leq\gamma]}$
$\leq$ $\sqrt{\Pr\big{[}|\mathbf{e}_{1},\zeta_{0}-\zeta_{0}^{\prime}|\leq\mu\gamma\mid\|\zeta_{0}\|\leq\gamma,\|\zeta_{0}^{\prime}\|\leq\gamma\big{]}}$
$\leq$ $\iota/2.$
Hence, we show that
$\Pr[F(y)-F(x_{0})\leq-\Phi]\geq 1-\iota.$
Appendix B Proof of Subsection 3.2
B.1 Proof of Lemma 3.11
Proof.
We only prove the first two items, as (3) and (4) follow from similar arguments. For each data $z\sim\mathcal{P}$, we know
@@
Then the conclusion (1) follows from Lemma 2.4.
@@
B.2 Proof of Lemma 3.12
@@
That is $\Delta_{i}-(\nabla F(x_{i})-\nabla F(x_{i-1}))$ is zero-mean and $\mathrm{nSG}(M\|x_{i}-x_{i-1}\|\sqrt{\log(dT/\iota)}/\sqrt{b_{i}})$ by applying Lemma 3.11. Then Equation (2) follows from applying Lemma 2.4.
The case for Hessian estimation involves some truncation. By Lemma 3.11, we know with probability at least $1-T\iota$, we have
$\Pr[\|H_{\tau_{t}}-\nabla^{2}F(x_{\tau_{t}})\|_{2}\geq\zeta_{3}]\leq\iota/T,$
with $\zeta_{3}=O(M\sqrt{\log(Td/\iota)/b})$. The similar concentration holds for $\Delta_{t}^{H}$. We truncate the distribution of $H_{\tau_{t}}$ around $\nabla^{2}F(x_{\tau_{t}})$, that is $\bar{H}_{\tau_{t}}=H_{\tau_{t}}\cdot\mathbf{1}\Big{(}\|H_{\tau_{t}}-\nabla^{2}F(x_{\tau_{t}})\|\leq\zeta_{3}\Big{)}$. It is straightforward to see that
$\|\operatorname*{\mathbb{E}}\bar{H}_{\tau_{t}}-\nabla^{2}F(x_{\tau_{t}})\|\leq\int_{\zeta_{3}}^{\infty}\Pr[\|H_{\tau_{t}}-\nabla^{2}F(x_{\tau_{t}})\|\geq\zeta]\mathrm{d}\zeta\leq\zeta_{3}/T.$
Truncate $\Delta_{t}^{H}$ in a similar way. Then by Lemma 3.11 we can show that
$\Pr[\|\bar{H}_{\tau_{t}}+\sum_{i=\tau_{t}+1}^{t}\overline{\Delta}_{i}^{H}-\operatorname*{\mathbb{E}}[\bar{H}_{\tau_{t}}+\sum_{i=\tau_{t}+1}^{t}\bar{\Delta}_{i}^{H}]\|^{2}\geq\frac{C}{2}(\frac{M^{2}}{b}+\sum_{i=\tau_{t}+1}^{t}\rho^{2}\|x_{i}-x_{i-1}\|^{2}/b_{i})\log^{2}(Td/\iota)]\leq\iota.$ (6)
Note that
$\|\operatorname*{\mathbb{E}}[\bar{H}_{\tau_{t}}+\sum_{i=\tau_{t}+1}^{t}\bar{\Delta}_{i}^{H}]-\nabla^{2}F(x_{t})\|=$ $\|\operatorname*{\mathbb{E}}[\bar{H}_{\tau_{t}}+\sum_{i=\tau_{t}+1}^{t}\bar{\Delta}_{i}^{H}]-\operatorname*{\mathbb{E}}[H_{\tau_{t}}+\sum_{i=\tau_{t}+1}^{t}\Delta_{i}^{H}]\|$ (7)
$\leq$ $\frac{C\log(Td/\iota)}{2T}(\frac{M}{\sqrt{b}}+\sum_{i=\tau_{t}+1}^{t}\rho\|x_{i}-x_{i-1}\|/\sqrt{b_{i}}).$
By union bound, we have
$\Pr[\|\bar{H}_{\tau_{t}}+\sum_{i=\tau_{t}+1}^{t}\overline{\Delta}_{i}^{H}\neq H_{t}]\leq T\iota/2.$ (8)
Equations (6),(7) and (8) complete the proof. ∎
B.3 Proof of Lemma 3.13
@@
We use the tree mechanism to privatize the gradients if no potential saddle point is met, and we use the Gaussian mechanism during escaping from the saddle point.
We first show the indistinguishability of the gradients. It suffices to consider the sensitivity of the gradient oracles.
@@
The privacy guarantee of gradients follows from the tree mechanism (Theorem 2.2), which means {~$\{\widetilde{\nabla}_{t}\}_{t}\approx_{\varepsilon/2,\delta/2}\{\widetilde{\nabla}_{t}^{\prime}\}_{t}$.
As for the Gaussian mechanism, note that for neighboring datasets with Hessian estimates $H_{t}$ and $H_{t}^{\prime}$ respectively, we know that
$\|H_{t}-H_{t}^{\prime}\|_{2}\leq\frac{M}{b}+\max_{t}\frac{\rho\|x_{t}-x_{t-1}\|}{b_{t}}.$
Note that we force a fresh estimate from $\mathcal{O}_{3}$ after escaping the saddle points for $\tau$ times, and in each time, we ensure that $\|x_{t}-x_{0}\|\leq\Xi$ and $H_{t}$ are used at most $\Gamma$ steps, which means the difference is
$\|(H_{t}-H_{t}^{\prime})(x_{t}-x_{0})\|_{2}\leq\|H_{t}-H_{t}^{\prime}\|\cdot\|x_{t}-x_{0}\|\leq(\frac{M}{b}+\max_{t}\frac{\rho\|x_{t}-x_{t-1}\|}{b_{t}})\cdot\Xi.$
Then the property of Gaussian Mechanism and composition finishes the privacy guarantee proof. ∎
B.4 Proof of Lemma 3.14
@@
By our setting of parameters, we know $\Xi\sqrt{\Gamma\tau}\leq\alpha/\rho$ and hence
$(\frac{G}{b}+\max_{t}\frac{\|g_{t}\|}{b_{t}})\leq 2\varepsilon\alpha/\sqrt{d},$
$(\frac{M}{b}+\max_{t}\frac{\rho\|x_{t}-x_{t-1}\|}{b_{t}})\cdot(2\Xi\sqrt{\Gamma\tau})\leq 2\varepsilon\alpha/\sqrt{d},$
Then our choice of $\sigma_{\mathrm{tree}}$ and $\sigma_{t}$ ensures the privacy guarantee by Lemma 3.13.
For any $t\in[T]$, if it is in the process of escaping from the saddle point, and $g_{t}=g+H(x_{t}-x_{\mathrm{anchor}})+\xi_{t}$, then Lemma 3.5 and our parameter settings ensure the accuracy on $g_{t}$. Consider the other case when $g_{t}=\widetilde{\nabla}_{t}$, we have
@@
$\|\nabla_{t}-\nabla F(x_{t})\|^{2}\lesssim$ $(\alpha^{2}+\sum_{i=\tau_{t}+1}^{t}\|g_{t}\|^{2}/b_{t})\log(nd/\iota)$
$\lesssim$ $(\alpha^{2}+\sum_{i=\tau_{t}+1}^{t}\|g_{t}\|\cdot\min\{\frac{\alpha\varepsilon}{\sqrt{d}},\alpha^{2}/\kappa\})\log(nd/\iota)$
@@
Hence we conclude that $(2)\lesssim\alpha\sqrt{\log(nd/\iota)}$.
Similarly, by Lemma 3.12, we can conclude that
$\|H_{t}-\nabla^{2}F(x_{t})\|^{2}\lesssim$ $(\rho\alpha+\sum_{i=\tau_{t}+1}\rho^{2}\|g_{t}\|^{2}/M^{2}b_{t})\log^{2}(nd/\iota)$
$\lesssim$ $(\rho\alpha+\sum_{i=\tau_{t}+1}^{t}\rho^{2}\|g_{t}\|\cdot(\alpha/\rho\kappa))\log^{2}(nd/\iota)$
$\lesssim$ $\rho\alpha\log^{2}(nd/\iota),$
which completes the proof. ∎
B.5 Proof of Lemma 3.15